g(x)=2x^2-5 g(3)
Do you square the 3 and then multiply by 2?

Answer by jim_thompson5910(35256)

Yes, you are correct.


g(x)=2x^2-5

g(3)=2(3)^2-5

g(3)=2(9)-5

g(3)=18-5

g(3)=13